# Bilinear spherical maximal function with fractal dilations

### v1

@@ Abstract.
In this paper, we investigate $L^{p}$-boundedness of the bilinear spherical maximal function associated with a general dilation set $E\subset{\mathbb{R}}_{+}$. We quantify the range of $L^{p}$-boundedness in terms of a dilation-invariant notion of upper Minkowski dimension of the set $E$. A particular case of this study, settles an open question of $L^{p}$-boundedness of the lacunary bilinear spherical maximal function at borderline cases $p_{1}=1$ or $p_{2}=1$ in dimension $d\geq 4$.
@@ 1. Introduction
Let $A_{t}f(x)$ denote the average of $f$ on the sphere of radius $t>0$ centered at $x\in{\mathbb{R}}^{d},~d\geq 2$. It is defined by
$A_{t}f(x)=\int_{\mathbb{S}^{d-1}}f(x-ty)d\sigma(y),$
where $d\sigma$ denotes the normalized surface measure on $\mathbb{S}^{d-1}.$ For a dilation set $E\subset{\mathbb{R}}_{+}$, the corresponding spherical maximal operator is defined by
@@
The $L^{p}$-bounds for the maximal operator $M_{E}$ have been extensively studied in the literature. Stein [SphStein] used Fourier analytic methods to obtain optimal $L^{p}$-bounds for the maximal function $M_{{\mathbb{R}}_{+}}f$ for $d\geq 3.$ The Fourier analytic approach does not extend to resolve the problem in dimension $d=2$. Later, Bourgain [CirBourgain] introduced some geometric agruments to settle the case of dimension $d=2$. We refer to the maximal function $M_{{\mathbb{R}}_{+}}f$ as the full spherical maximal function and denote it by $M_{full}f$. At the end-point $p=\frac{d}{d-1}$, Bourgain [sphBourgain] proved the restricted weak-type bounds in dimensions $d\geq 3$, while it was shown by Seeger, Tao and Wright [STW] that the restricted weak-type inequality fails in dimension $d=2$.
Under restrictions on the dilation set $E\subset{\mathbb{R}}_{+}$, the $L^{p}$-boundedness of $M_{E}$ extends to a larger range of $L^{p}$-spaces. The range of $p$ depends on the geometry and appropriate notion of size of the set $E$. For a general dilation set $E\subset{\mathbb{R}}_{+}$, Seeger, Wainger and Wright [SWW1] obtained sharp $L^{p}$-bounds for $p>p(E)$, where $p(E)$ depends on the dilation-invariant notion of the Minkowski dimension. The end-point cases were resolved in [STW]. A particularly interesting case arises when we consider $E$ to be a lacunary set. In this case, the $L^{p}$-bounds of the operator $M_{E}$ (denoted by $M_{lac}$), holds for all $p>1$, see [lacsphCalderon, CoifmanWeiss, RubioDuoandi]. It is conjectured that the lacunary maximal function $M_{lac}$ is of weak-type $(1,1)$, however the problem remains open till date. We refer to  [CladekKrause] for recent developments in this direction. The sharp range of sparse bounds for $M_{full}$ and $M_{lac}$ was obtained by Lacey [sparseLacey] using continuity estimates and $L^{p}$-improving bounds.
When $E\subset[1,2]$, the corresponding maximal operator $M_{E}$ is referred to as the local maximal operator. Due to the local nature of the operator, it satisfies improving $L^{p}\to L^{q}$-bounds for some $q\geq p$. Such estimates for the local operator $M_{E}$, in terms of the upper Minkowski dimension and the Assouad spectrum of $E$, have been studied in [AHRS, RoosSeeger, LRZZ]. In these results, the Fourier decay of surface measures plays a key role, and the fractal geometry of $E$ is captured through the covering numbers and Minkowski dimension.
Let us recall the definition of dilation-invariant notion of the Minkowski dimension, see [SWW1] for more details.
@@ 1.1. Bilinear spherical maximal function
The bilinear spherical maximal function associated to a given dilation set $E\subset{\mathbb{R}}_{+}$ is defined by
@@
where the bilinear spherical average $\mathcal{A}_{t}(f_{1},f_{2})(x)$ is defined by
@@
The bilinear spherical averages first appeared in the work [GGIPS]. The $L^{p}$-estimates for the full maximal operator $\mathcal{M}_{full}:=\mathcal{M}_{{\mathbb{R}}_{+}}$ were studied in [BGHHO, GHH, HHY]. The sharp range of $L^{p}$-boundedness, including the end-point cases, were obtained by Jeong and Lee [JeongLee] for dimensions $d\geq 2$. They introduced the method of slicing and gave a simple proof of optimal $L^{p}$-estimates for the operator $\mathcal{M}_{full}$. The slicing argument provides with a decomposition of $\mathbb{S}^{2d-1}$ into a family of lower dimensional spheres. More precisely, Jeong and Lee [JeongLee] showed that if $F$ is a continuous function defined on ${\mathbb{R}}^{2d},d\geq 2$ and $(y,z)\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$, then
@@
The formula as above yields the following pointwise estimate.
@@
where $M_{HL}f$ is the Hardy-Littlewood maximal function. Using the symmetry of the bilinear spherical maximal function in $f$ and $g$, the boundedness region for the operator $\mathcal{M}_{full}$ is determined by using $L^{p}$-estimates of Hardy-Littlewood maximal function $M_{HL}f(x)$ and spherical maximal function $M_{full}g(x)$ for $d\geq 2.$ In dimension $d=1$, the optimal range of boundedness of $\mathcal{M}_{full}$ was obtained in [ChristZhou, DosidisRamos, BCSS].
Jeong and Lee [JeongLee] also obtained $L^{p}$-improving estimates for the local version of bilinear spherical maximal function using the slicing method. Later, in [BCSS], the authors improved the range of $L^{p}$-improving estimates for the local bilinear spherical maximal function. Observe that for a general dilation set $E$, the slicing argument always leads to pointwise domination of the bilinear maximal function $\mathcal{M}_{E}(f,g)$ by a product of the Hardy-Littlewood maximal function and the full spherical maximal function as in the estimate (1.2). This does not improve the range of $L^{p}$-boundedness of the operator $\mathcal{M}_{E}$. Therefore, the slicing argument has its limitations in dealing with the bilinear maximal function $\mathcal{M}_{E}(f,g)$ associated with a dilation set $E$.
The $L^{p}$-estimates for the bilinear lacunary spherical maximal operator $\mathcal{M}_{lac}=:\mathcal{M}_{E}$, where $E$ is a lacunary set, were obtained by Borges and Foster [BorgesFoster] for dimensions $d\geq 2$, and by Christ and Zhou  [ChristZhou] for dimension $d=1$. The proof for dimensions $d\geq 2$ relies mainly on an $L^{2}\times L^{2}\rightarrow L^{1}$-decay estimate in the frequency parameter for single scale average $\mathcal{A}_{1}(f,g)$, where functions $f$ or $g$ are frequency localized in an appropriate annular region formed by Littlewood-Paley decomposition. Then, the desired estimates for the bilinear lacunary maximal function are obtained by a boot-strapping type argument. The decay estimate at the point $(2,2,1)$ is obtained by a combination of slicing argument (1.1) and Fourier decay property of $\widehat{d\sigma}$. Due to unavailability of the slicing method and lack of appropriate Fourier decay of $\widehat{d\sigma}$ in dimension $d=1$, it is significantly more difficult to obtain $L^{p}$-estimates for the bilinear lacunary maximal function in dimension $d=1$. Christ and Zhou  [ChristZhou] overcame these issues and proved trilinear smoothing estimates for bilinear averaging operators along certain curves, including the averages along the circle. The trilinear smoothing estimates play a key role in obtaining decay estimates for frequency localized pieces and finally obtaining the desired $L^{p}$-estimates for $\mathcal{M}_{lac}$. Here we summarize the results for the bilinear lacunary maximal operator $\mathcal{M}_{lac}$.
@@
[ChristZhou, BorgesFoster] Let $1\leq p_{1},p_{2}\leq\infty$ with $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}$. Then, $\mathcal{M}_{lac}$ maps $L^{p_{1}}(\mathbb{R}^{d})\times L^{p_{2}}(\mathbb{R}^{d})$ to $L^{p}(\mathbb{R}^{d})$ if
@@
Moreover, the weak-type bounds $\mathcal{M}_{lac}:L^{1}(\mathbb{R}^{d})\times L^{\infty}(\mathbb{R}^{d})\rightarrow L^{1,\infty}(\mathbb{R}^{d})$ and $\mathcal{M}_{lac}:L^{\infty}(\mathbb{R}^{d})\times L^{1}(\mathbb{R}^{d})\rightarrow L^{1,\infty}(\mathbb{R}^{d})$ hold.
@@
Christ and Zhou [ChristZhou] showed that the strong-type $L^{p}$-bounds for $\mathcal{M}_{lac}$ do not hold if either $p_{1}=1$ or $p_{2}=1$. However, in higher dimensions, $L^{p}$-boundedness of $\mathcal{M}_{lac}$ is known to hold for the range $p_{1}=1$ and $p_{2}\in(\frac{d}{d-1},\infty)$, or $p_{2}=1$ and $p_{1}\in(\frac{d}{d-1},\infty)$. The question concerning the remaining points on the line segments $p_{1}=1$ and $p_{2}=1$ remained unresolved. In this paper, we address this problem and extend the range of $L^{p}$-boundedness of $\mathcal{M}_{lac}$ to the borderline cases in $d\geq 4$ excepting the end-point point $(1,1)$. Indeed, we obtain new results for bilinear spherical maximal function associated with general dilation sets. The study of local bilinear spherical maximal function with a general dilation set $E\subset[1,2]$ was carried out by Borges, Foster and Ou [BFO]. They proved Sobolev smoothing estimates at the point $(2,2,2)$ for the local maximal operator along hypersurfaces that exhibit Fourier decay of order $\frac{1}{(1+|\xi|)^{\alpha}}$ with $\alpha>\frac{d}{2}$. They also obtained necessary conditions for the boundedness of $\mathcal{M}_{E}$ in terms of the upper Minkowski dimension of the underlying fractal dilation set. The Sobolev smoothing estimates give a simplified derivation of the continuity estimates for such local maximal operators associated with general dilation sets $E$. They also deduced $L^{p}$-improving bounds and sparse bounds for the associated bilinear maximal operators. However, when restricted to spherical maximal functions, this does not improve the $L^{p}$-boundedness range for $\mathcal{M}_{E}$ beyond that of $\mathcal{M}_{full}$.
The sparse bounds for the bilinear spherical maximal function $\mathcal{M}_{full}$ were studied by Borges et al. in [BFOPZ]. Extensions of these results to some general surfaces have also been explored; for instance, see [CGHHS, LS] for $L^{p}$-estimates and [RSS, PS] for sparse bounds. Additionally, the multilinear extension of the maximal function associated with spheres and other general surfaces has been studied in [Dosidis, SS, GHHP, CLS, Gao].
@@ 2. Main results
In this paper, we explore alternate of slicing method to obtain new results concerning $L^{p}$-boundedness of the bilinear maximal operator $\mathcal{M}_{E}$ associated with general dilation set $E$. We intend to quantify the range of exponents $(p_{1},p_{2},p)$ for the $L^{p}$-boundedness of the operator $\mathcal{M}_{E}$ in terms of upper Minkowski dimension of the set $E$. We exploit the Fourier decay property of the surface measure in a refined manner and perform a scale-adapted discretization of $E$ to capture the fractal geometry of the set $E$. These ideas, along with suitable adaptation of bilinear Calderón-Zygmund theory, yield new results for the maximal operator $\mathcal{M}_{E}$. In order to describe our results, we need the following notation.
Let $\Omega({P_{1},P_{2},\dots,P_{k}})$ denote the convex hull of points $\{P_{1},P_{2},\dots,P_{k}\}\subset{\mathbb{R}}^{2}.$ Let $d\geq 2$ and set $O=(0,0),\;A=(1,0),\;B=(0,1),\;P_{0}=\big(\frac{2d-1}{2d},\frac{2d-1}{2d}\big),\;Q_{0}=\big(1,\frac{d-1}{d}\big),\;\text{and}~R_{0}=\big(\frac{d-1}{d},1\big)$. For $0\leq\beta\leq 1$ consider the point
$P=P(d,\beta)=\begin{cases}\left(\frac{2d-2}{2d-2+\beta},\frac{2d-2}{2d-2+\beta}\right)&\beta<\frac{2d-2}{2d-1},\\ P_{0}&\beta\geq\frac{2d-2}{2d-1}.\end{cases}$
Further, for $d\geq 4$, we define
$Q=Q(d,\beta)=\begin{cases}\left(1,\frac{d-3}{d-3+\beta}\right)&\beta<\frac{d-3}{d-1},\\ Q_{0}&\beta\geq\frac{d-3}{d-1},\end{cases}\qquad\text{and}\qquad R=R(d,\beta)=\begin{cases}\left(\frac{d-3}{d-3+\beta},1\right)&\beta<\frac{d-3}{d-1},\\ R_{0}&\beta\geq\frac{d-3}{d-1}.\end{cases}$
Throughout this paper, we shall assume that exponents $(p_{1},p_{2},p)$ satisfying $1\leq p_{1},p_{2}\leq\infty$ and the Hölder condition $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}$ unless specified otherwise. We also use the notation $A\lesssim B$ to denote that $A\leq CB$ for some $C>0$.
The following is the main result of the paper.
@@
    $d=2,3$, and $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(\{O,A,Q_{0},P,R_{0},B\})\setminus\{A,B,PQ_{0},PR_{0}\}$.
@@
    $d\geq 4$, and $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(\{O,A,Q,P,R,B\})\setminus\{A,B,PQ,PR\}$.
Moreover, if the set $E\subseteq[1,2]$ has upper Minkowski dimension $\beta$, then the local bilinear maximal function $\mathcal{M}_{E}$ satisfies restricted weak-type estimate $L^{p_{1},1}(\mathbb{R}^{d})\times L^{p_{2},1}(\mathbb{R}^{d})\rightarrow L^{p,\infty}(\mathbb{R}^{d})$ whenever $\Big(\frac{1}{p_{1}},\frac{1}{p_{2}}\Big)$ lies on the line segments $QP$ or $PR$.
@@
The lacunary bilinear maximal function corresponds to the case $\beta=0$ in the theorem above. Therefore, we have $L^{p}$-boundedness of the lacunary bilinear spherical maximal operator $\mathcal{M}_{lac}$ in the borderline cases in dimension $d\geq 4.$
@@
2.1. Strategy of proof
As indicated earlier, the slicing method does not improve the range of $L^{p}$-estimates for the bilinear spherical maximal functions associated with general dilation sets beyond that of the full maximal operator $\mathcal{M}_{full}$. We follow the following strategy to prove Theorem 2.1.
  • First, we perform Littlewood-Paley decomposition of both the functions in frequency space. The $L^{p}$-estimates for the maximal function acting on low frequency terms and terms with mixed type frequencies can be dealt with directly. In this case, we could dominate the maximal function by a product of Hardy-Littlewood maximal functions, see Lemma 3.1. This step reduces the problem to proving the desired bounds for terms with high frequencies.
  • Next, we consider the local maximal operator associated with the set $E_{0}=E\cap[1,2]$. In Section 4 we carry out the decomposition of multiplier as used by Heo, Hong and Yang  [HHY].
  • We derive
@@ 2.1. Strategy of proof
    estimates, see Lemma 5.2 with suitable decay for the operator norm depending on the frequency parameter and the upper Minkowski content of $E_{0}$. The second estimate that we obtain is an end-point estimate for the operator. We use an appropriate space localization of the operator and establish end-point $L^{1}\times L^{1}\to L^{1/2}$-estimate, see Lemma 5.3. This end-point estimate has growth in terms of the frequency parameter and the number of intervals used in discretization of the $E_{0}$.
  • The interpolation argument between the growth estimate at end-point and the decay estimate yields a region of $L^{p}$-boundedness of $\mathcal{M}_{E_{0}}$ depending on upper Minkowski dimension of $E_{0}$.
  • Finally, in Section 6, we extend the results obtained for local maximal operator to the global operators associated with $E\subset{\mathbb{R}}_{+}$. This part is done by using appropriate rescaling arguments and adaptation of the bilinear Calderón-Zygmund theory to the underlying situation. We exploit the methods used in [ChristZhou] and [BCSS2].

### v2

@@ Abstract.
In this paper, we investigate $L^{p}$-boundedness of the bilinear spherical maximal function associated with a general set $E\subset{\mathbb{R}}_{+}$. We quantify the range of $L^{p}$-boundedness in terms of a dilation-invariant notion of upper Minkowski dimension of the set $E$. A particular case of this study, settles an open question of $L^{p}$-boundedness of the lacunary bilinear spherical maximal function at borderline cases $p_{1}=1$ or $p_{2}=1$ in dimension $d\geq 3$.
@@ 1. Introduction
The spherical average of $f:{\mathbb{R}}^{d}\rightarrow{\mathbb{C}},~d\geq 2,$ is defined by
$A_{t}f(x)=\int_{\mathbb{S}^{d-1}}f(x-ty)d\sigma(y),~t>0,$
where $d\sigma$ denotes the normalized surface measure on the sphere $\mathbb{S}^{d-1}.$ Given a set $E\subset{\mathbb{R}}_{+}$, the corresponding spherical maximal operator is defined by
@@
Stein [38] used Fourier analytic methods to prove that the maximal operator $M_{{\mathbb{R}}_{+}}f$ is bounded in $L^{p}$ for $p>\frac{d}{d-1}$ for $d\geq 3$. This range of $p$ is sharp upto up to the end-point. The maximal operator $M_{{\mathbb{R}}_{+}}f$ is referred to as the full spherical maximal operator and it is denoted by $M_{full}f$. The Fourier analytic approach has a limitation in resolving the problem in dimension $d=2$. Bourgain [9] introduced a geometric method to settle the case of dimension $d=2$. He proved that $M_{full}f$ is bounded in $L^{p}$ for $p>2$ for $d=2$. At the end-point $p=\frac{d}{d-1}$, Bourgain [8] proved restricted weak-type bounds in dimensions $d\geq 3$, while it was shown by Seeger, Tao and Wright [35] that the restricted weak-type inequality at the end-point $p=\frac{d}{d-1}$ fails in dimension $d=2$.
Under suitable restrictions on the set $E\subset{\mathbb{R}}_{+}$, the $L^{p}$-boundedness of $M_{E}$ extends to a larger range of $L^{p}$-spaces. The range of $p$ depends on the geometry and appropriate notion of size of the set $E$. For a general set $E\subset{\mathbb{R}}_{+}$, Seeger, Wainger and Wright [36] obtained sharp $L^{p}$-bounds for $p>p(E)$, where $p(E)$ depends on the dilation-invariant notion of the Minkowski dimension. The end-point cases were resolved in [35]. A particularly interesting case arises when we consider $E$ to be a lacunary set. In this case, the $L^{p}$-bounds of the operator $M_{E}$ (denoted by $M_{lac}$), holds for all $p>1$, see [10, 15, 18]. It is conjectured that the lacunary maximal operator $M_{lac}$ is of weak-type $(1,1)$, however the problem remains open till date. We refer to  [14] for recent developments in this direction. Further, we would like to refer the reader to Lacey [27] for the sharp range of sparse domination results for the maximal operators $M_{full}$ and $M_{lac}$. The continuity estimates and $L^{p}$-improving bounds for local variants of spherical maximal function play key roles in proving sparse domination results.
When $E\subset[1,2]$, the corresponding maximal operator $M_{E}$ is referred to as the local maximal operator. Due to the local nature of the operator, it satisfies improving $L^{p}\to L^{q}$-bounds for a range of $p$ and $q$ with $q\geq p$. Such estimates for the local operator $M_{E}$, in terms of the upper Minkowski dimension and the Assouad spectrum of $E$, have been studied in [1, 34, 28]. In these results, the Fourier decay of surface measures plays a key role, and the fractal geometry of $E$ is captured through the covering numbers and Minkowski dimension.
Let us recall the definition of dilation-invariant notion of the Minkowski dimension, see [36] for more details.
@@ 1.1. Bilinear spherical maximal function
The bilinear spherical maximal function associated to a given set $E\subset{\mathbb{R}}_{+}$ is defined by
@@
where the bilinear spherical average $\mathcal{A}_{t}(f_{1},f_{2})(x)$ is given by
@@
The bilinear spherical averages first appeared in [20]. The $L^{p}$-estimates for the full maximal operator $\mathcal{M}_{full}:=\mathcal{M}_{{\mathbb{R}}_{+}}$ were studied in [2, 22, 24]. The sharp range of $L^{p}$-boundedness, including the end-point cases, were obtained by Jeong and Lee [26] for dimensions $d\geq 2$. They introduced the method of slicing and gave a simple proof of optimal $L^{p}$-estimates for the operator $\mathcal{M}_{full}$. The slicing argument provides with a decomposition of $\mathbb{S}^{2d-1}$ into a family of lower dimensional spheres. More precisely, Jeong and Lee [26] showed that if $F$ is a continuous function defined on ${\mathbb{R}}^{2d},d\geq 2$ and $(y,z)\in{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}$, then
@@
The formula as above yields the following pointwise estimate
@@
where $M_{HL}f$ is the Hardy-Littlewood maximal function. Using the symmetry of the bilinear spherical maximal function in $f$ and $g$, the boundedness region for the operator $\mathcal{M}_{full}$ is determined by using $L^{p}$-estimates of Hardy-Littlewood maximal function $M_{HL}f(x)$ and spherical maximal function $M_{full}g(x)$ for $d\geq 2.$ In dimension $d=1$, the optimal range of boundedness of $\mathcal{M}_{full}$ was obtained in [13, 16, 3].
Jeong and Lee [26] also obtained $L^{p}$-improving estimates for the local version of bilinear spherical maximal function using the slicing method. Later, in [3], the authors improved the range of $L^{p}$-improving estimates for the local bilinear spherical maximal function. Observe that for a general set $E$, the slicing argument always leads to pointwise domination of the bilinear maximal function $\mathcal{M}_{E}(f,g)$ by a product of the Hardy-Littlewood maximal function and the full spherical maximal function as in the estimate (1.2). This does not improve the range of $L^{p}$-boundedness of the operator $\mathcal{M}_{E}$. Therefore, the slicing argument has its limitations in dealing with the bilinear maximal function $\mathcal{M}_{E}(f,g)$ associated with a set $E$.
The $L^{p}$-estimates for the bilinear lacunary spherical maximal operator $\mathcal{M}_{lac}=:\mathcal{M}_{E}$, where $E$ is a lacunary set, were obtained by Borges and Foster [7] for dimensions $d\geq 2$, and by Christ and Zhou  [13] for dimension $d=1$. The proof for dimensions $d\geq 2$ relies mainly on an $L^{2}\times L^{2}\rightarrow L^{1}$-decay estimate in the frequency parameter for single scale average $\mathcal{A}_{1}(f,g)$, where either $f$ or $g$ is frequency localized in an appropriate annular region formed by Littlewood-Paley decomposition. Then, the desired estimates for the bilinear lacunary maximal function are obtained by a boot-strapping type argument. The decay estimate at the point $(2,2,1)$ is obtained by a combination of slicing argument (1.1) and Fourier decay property of $\widehat{d\sigma}$. Due to unavailability of the slicing method and lack of appropriate Fourier decay of $\widehat{d\sigma}$ in dimension $d=1$, it is significantly more difficult to obtain $L^{p}$-estimates for the bilinear lacunary maximal function in dimension $d=1$. Christ and Zhou  [13] overcame these issues and proved trilinear smoothing estimates for bilinear averaging operators along certain curves, including the averages along the circle. The trilinear smoothing estimates play a key role in proving decay estimates at the point $(2,2,1)$ for frequency localized pieces and finally in obtaining the desired $L^{p}$-estimates for $\mathcal{M}_{lac}$. Here we summarize the results for the bilinear lacunary maximal operator $\mathcal{M}_{lac}$.
@@
[13, 7] Let $1\leq p_{1},p_{2}\leq\infty$ with $\frac{1}{p}=\frac{1}{p_{1}}+\frac{1}{p_{2}}$. Then, $\mathcal{M}_{lac}$ maps $L^{p_{1}}(\mathbb{R}^{d})\times L^{p_{2}}(\mathbb{R}^{d})$ to $L^{p}(\mathbb{R}^{d})$ if
@@
Moreover, the weak-type boundedness $\mathcal{M}_{lac}:L^{1}(\mathbb{R}^{d})\times L^{\infty}(\mathbb{R}^{d})\rightarrow L^{1,\infty}(\mathbb{R}^{d})$ and $\mathcal{M}_{lac}:L^{\infty}(\mathbb{R}^{d})\times L^{1}(\mathbb{R}^{d})\rightarrow L^{1,\infty}(\mathbb{R}^{d})$ hold.
@@
Christ and Zhou [13] showed that the strong-type $L^{p}$-bounds for $\mathcal{M}_{lac}$ do not hold if either $p_{1}=1$ or $p_{2}=1$. However, in higher dimensions, $L^{p}$-boundedness of $\mathcal{M}_{lac}$ is known to hold for the range $p_{1}=1$ and $p_{2}\in(\frac{d}{d-1},\infty)$, or $p_{2}=1$ and $p_{1}\in(\frac{d}{d-1},\infty)$. The question concerning the remaining points on the line segments $p_{1}=1$ and $p_{2}=1$ remained unresolved. In this paper, we address this problem and extend the range of $L^{p}$-boundedness of $\mathcal{M}_{lac}$ to the borderline cases in $d\geq 4$ excepting the end-point point $(1,1)$. Indeed, we obtain new results for bilinear spherical maximal function associated with general sets. The study of local bilinear spherical maximal function with a general set $E\subset[1,2]$ was carried out by Borges, Foster and Ou [6]. They proved Sobolev smoothing estimates at the point $(2,2,2)$ for the local maximal operator along hypersurfaces that exhibit Fourier decay of order $\frac{1}{(1+|\xi|)^{\alpha}}$ with $\alpha>\frac{d}{2}$. They also obtained necessary conditions for the boundedness of $\mathcal{M}_{E}$ in terms of the upper Minkowski dimension of the underlying fractal set. The Sobolev smoothing estimates give a simplified derivation of the continuity estimates for such local maximal operators associated with general sets $E$. They also deduced $L^{p}$-improving bounds and sparse bounds for the associated bilinear maximal operators. However, when restricted to spherical maximal functions, this does not improve the $L^{p}$-boundedness range for $\mathcal{M}_{E}$ beyond that of $\mathcal{M}_{full}$.
The sparse bounds for the bilinear spherical maximal function $\mathcal{M}_{full}$ were studied by Borges et al. in [5]. Extensions of these results to some general surfaces have also been explored; for instance, see [11, 29] for $L^{p}$-estimates and [33, 32] for sparse bounds. Additionally, the multilinear extension of the maximal function associated with spheres and other general surfaces has been studied in [17, 37, 21, 12, 19].
@@ 2. Main results
In this paper, we explore and develop an alternative approach to the slicing method in order to obtain new results concerning $L^{p}$-boundedness of the bilinear maximal operator $\mathcal{M}_{E}$ associated with general set $E$. We intend to quantify the range of exponents $(p_{1},p_{2},p)$ for the $L^{p_{1}}\times L^{p_{2}}\rightarrow L^{p}$-boundedness of the operator $\mathcal{M}_{E}$ in terms of upper Minkowski dimension of the set $E$. We exploit the Fourier decay property of the surface measure in a refined manner and perform a scale-adapted discretization of the set $E$ to capture fractal geometry of the underlying set. These ideas, along with suitable adaptation of bilinear Calderón-Zygmund theory, yield new results for the maximal operator $\mathcal{M}_{E}$. In order to describe our results, we need the following notation.
Let $\Omega({P_{1},P_{2},\dots,P_{k}})$ denote the convex hull of points $P_{1},P_{2},\dots,P_{k}\in{\mathbb{R}}^{2}.$ Let $d\geq 2$ and set $O=(0,0),\;A=(1,0),\;B=(0,1),\;P_{0}=\big(\frac{2d-1}{2d},\frac{2d-1}{2d}\big),\;Q_{0}=\big(1,\frac{d-1}{d}\big),\;\text{and}~R_{0}=\big(\frac{d-1}{d},1\big)$. For $0\leq\beta\leq 1$ consider the point
$P=P(d,\beta)=\begin{cases}\left(\frac{2d-1-\beta}{2d-1},\frac{2d-1-\beta}{2d-1}\right),&0\leq\beta<\frac{2d-1}{2d},\\ P_{0},&\frac{2d-1}{2d}\leq\beta\leq 1.\end{cases}$
Further, for $d\geq 3$, we define $Q=Q(d,\beta)$ and $R=R(d,\beta)$ by
[…]
By a direct calculation, one can observe that for $d\leq 5$ we have that $\frac{d-3}{d-1}\leq\frac{d-2}{d+1}\leq\frac{1}{2}$. Therefore, for $d\leq 5$, the points $Q$ and $R$ are given by
[…]
If $d\geq 6$, then $\frac{1}{2}<\frac{d-2}{d+1}<\frac{d-3}{d-1}.$ Consequently, the points $Q$ and $R$ are
[…]
The following are the main results of the paper.
@@
    $d=2$, and $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(O,A,Q_{0},P,R_{0},B)\setminus\{A,B,PQ_{0},PR_{0}\}$.
@@
    $d\geq 3$, and $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(O,A,Q,P,R,B)\setminus\{A,B,PQ,PR\}$.
@@
Theorem 2.2.
Let $d=1$ and $0<p_{1},p_{2},p\leq\infty$ with $\frac{1}{p_{1}}+\frac{1}{p_{2}}=\frac{1}{p}$. Assume that $E\subset{\mathbb{R}}_{+}$ has an upper Minkowski dimension $\beta$ and define
$D_{1}=\begin{cases}\left(1,0\right)&\beta<\frac{1}{2},\\ \left(\frac{1}{2\beta},0\right)&\beta\geq\frac{1}{2},\end{cases}\qquad D_{2}=\begin{cases}\left(1-\beta,1-\beta\right)&\beta<\frac{1}{2},\\ \left(\frac{1}{2},\frac{1}{2}\right)&\beta\geq\frac{1}{2},\end{cases}\quad\text{and}\quad D_{3}=\begin{cases}\left(0,1\right)&\beta<\frac{1}{2},\\ \left(0,\frac{1}{2\beta}\right)&\beta\geq\frac{1}{2}.\end{cases}$
Then, the operator $\mathcal{M}_{E}$ is bounded from $L^{p_{1}}(\mathbb{R})\times L^{p_{2}}(\mathbb{R})$ into $L^{p}(\mathbb{R})$ when $(\frac{1}{p_{1}},\frac{1}{p_{2}})\in\Omega(O,D_{1},D_{2},D_{3})\setminus\{D_{1}D_{2},D_{2}D_{3}\}$.
[…]
Figure 3. Boundedness regions for the operator $\mathcal{M}_{E}$ in dimension $d=1$.
Note that the bilinear lacunary spherical maximal function corresponds to the case $\beta=0$, which was addressed by Christ and Zhou [13] for dimension $d=1$ using trilinear smoothing inequalities for high frequency terms. We recover their result as a special case of Theorem 2.2. Further, as a consequence of Theorem 2.1, we obtain $L^{p}$-boundedness of the lacunary bilinear spherical maximal operator $\mathcal{M}_{lac}$ in the borderline cases in dimension $d\geq 3.$
@@
Methodology of proofs
As indicated earlier, the slicing method does not improve the range of $L^{p}$-estimates for the bilinear spherical maximal functions associated with general sets beyond that of the full maximal operator $\mathcal{M}_{full}$. We overcome this difficulty by adopting to the following strategy to prove Theorem 2.1.
  • First, we perform Littlewood-Paley decomposition of both the functions in frequency space. The $L^{p}$-estimates for the maximal function acting on low frequency terms and terms with mixed type frequencies are dealt with directly. In this case, we could dominate the maximal function by a product of Hardy-Littlewood maximal functions, see Lemma 3.1. This step reduces the problem to proving the desired bounds for terms with high frequencies.
  • Next, we consider the local maximal operator associated with the set $E_{0}=E\cap[1,2]$ and carry out the decomposition of multiplier as used by Heo, Hong and Yang  [24]. We derive
@@
    estimates, see Lemma 5.1 and Lemma 5.2, with suitable decay for the operator norm depending on the frequency parameter and the upper Minkowski content of $E_{0}$. The second estimate that we obtain is an end-point estimate for the operator. We use an appropriate space localization of the operator and establish end-point $L^{1}\times L^{1}\to L^{1/2}$-estimate, see Lemma 5.3. This end-point estimate has growth in terms of the frequency parameter and the number of intervals used in discretization of the $E_{0}$.
  • The interpolation argument between the growth estimate at end-points and the decay estimate at points mentioned above yields a region of $L^{p}$-boundedness of $\mathcal{M}_{E_{0}}$ depending on upper Minkowski dimension of $E_{0}$.
  • Finally, we extend the results obtained for local maximal operator to the global operators associated with $E\subset{\mathbb{R}}_{+}$. This part is done by using appropriate rescaling arguments and adaptation of the bilinear Calderón-Zygmund theory to the underlying situation. We exploit the methods used in [13] and [4].
Organization of the paper
In Section 3, we decompose the operator using Littlewood-Paley decomposition of both the functions. In Section 4, we decompose the multiplier corresponding to the bilinear spherical average. In Section 5, we obtain $L^{p}$-estimates for local bilinear spherical maximal function. In Section 6 and Section 7, we prove the main results Theorem 2.1 and Theorem 2.2 respectively.